### 19 Mar 2026 v1

# Adaptive Layerwise Perturbation: Unifying Off-Policy Corrections for LLM RL

Abstract
Off-policy problems such as policy staleness and training–inference mismatch, has become a major bottleneck for training stability and further exploration for LLM RL. To enhance inference efficiency, the distribution gap between the inference and updated policy grows, leading to heavy-tailed importance ratios. Heavy-tailed ratios arise when the policy is locally sharp, which further inflates sharp gradients and can push updates outside the trust region. To address this, we propose Adaptive Layerwise Perturbation (ALP) by injecting small learnable perturbations into input hidden states of each layer during updates, which is used as the numerator of the importance ratio against the unchanged inference policy in the objective. Intuitively, by adding controlled noise to intermediate representations, ALP prevents the updated policy from deviating too sharply from the inference policy, and enlarges the policy family to cover the inference policy family with mismatch noises. Hence, the flattened distribution can naturally tighten the updated and inference policy gap and reduce the tail of importance ratios, thus maintaining training stability. This is further validated empirically. Experiments on single-turn math and multi-turn tool-integrated reasoning tasks show that ALP not only improves final performance, but also avoid blow up of importance ratio tail and KL spikes during iterative training, along with boosted exploration. Ablations show that representation-level perturbations across all layers are most effective, substantially outperforming partial-layer and logits-only variants.
Figure 1: Left: Visualization of Adaptive Layerwise Perturbation (ALP), where layerwise small perturbation variable is added to the model during training to cover the layerwise bias between training and inference policies. Right: the comparison between full training with perturbation (ALP) or without perturbation (Bypass) across $1840$ training steps. Without perturbation, the policy distribution becomes more spiky and brittle, leading to blow-up tails. In contrast, ALP smooths policies, tightens the envelope and stabilizes the importance ratio.
1 Introduction
The trade-off between training efficiency and stability is a central challenge in reinforcement learning (RL) for large language models (LLMs). To increase efficiency, practical systems often optimize a policy using rollouts generated by a different distribution, which is known as off-policy. Off-policy effects arise from multiple sources. First, the same batch of trajectories is used across several policy updates. Second, modern inference engines such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) introduce training–inference mismatch due to quantization, batching, and kernel-level differences, even when weights are nominally identical (Yao et al., 2025). In multi-turn agentic settings, this mismatch can be further amplified by inference distribution shifts induced by tool feedback and out-of-distribution observations (Jiacai et al., 2025). These factors lead to heavy-tailed importance ratios, KL spikes, and brittle optimization dynamics, making robust off-policy RL an urgent problem.
A prominent line of work stabilizes training by modifying importance ratios. One approach changes the aggregation way of importance ratios in the objective. It replaces the ratio at each token index in GRPO (Shao et al., 2024) with the multiplication of ratios in the whole trajectory (Zheng et al., 2025), but this does not consider training-inference mismatch. Another approach targets on training-inference mismatch. For example, Bypass is a baseline mentioned in (Jiacai et al., 2025; Yao et al., 2025) that uses the ratio of the updated and rollout policies. TIS/MIS (Yao et al., 2025; Jiacai et al., 2025) further correct mismatch by multiplying an auxiliary ratio of the proxy policy (an old policy acting as the anchor) and the rollout policy. Then, they truncate or mask tokens whose auxiliary ratio exceeds a threshold. Although effective at preventing catastrophic collapse, these methods introduce practical limitations: they split off-policy effects into two separately truncated ratios, which can over-truncate updates that are otherwise within a valid trust region, thereby inducing additional bias and leading to early plateau. We defer a detailed discussion of related work to Section 2.
We argue that off-policy instability in LLM RL is not solely a bookkeeping issue of “which importance ratio to use,” but also a geometry issue: the noisy update may push the policy toward sharp, brittle regions where small distribution shifts (from staleness or inference mismatch) cause large changes in action probabilities. This motivates a perturbation-sampling perspective from model smoothness and distributional learning approaches (Shen & Meinshausen, 2023; Hao et al., 2025): when the dominant failure mode is system-induced noise and staleness, a natural defense is fighting noise with structured noise.
Specifically, we propose Adaptive Layerwise Perturbation (ALP) by injecting a learnable Gaussian perturbation $\delta\sim\mathcal{N}(0,\sigma^{2}I)$  [^1] The perturbation distribution is not restricted to Gaussian, it only needs to have zero mean. into the input hidden states across layers during policy updates and producing a perturbed training distribution $\pi^{\mathrm{train}}_{\theta,\sigma}$ (Figure 1: Left). We then optimize the objective using a single unified ratio of $\pi^{\mathrm{train}}_{\theta,\sigma}$ and the unperturbed inference policy i.e., perturbation is applied only to the numerator during training.
Intuitively, ALP can smooth the local optimization landscape to suppress heavy-tailed ratio excursions, and enlarge the training distribution family. Moreover, this is testified by the full-training comparison in Figure 1 (Right). Starting from the same base model, the training policy without perturbation (Bypass) gradually gets sharp and brittle, and the importance ratio tail finally explode. In contrast, the training policy with perturbation (ALP) stays stable and controlling the mismatch between the training and inference system, thus tightening the ratio quantile envelop. Beyond stability, by enlarging effective support and preventing premature concentration on brittle modes, ALP also encourages exploration, particularly in multi-turn settings where compounding errors can reduce coverage.
Both our theoretical and empirical findings show that ALP significantly improves robustness and final performance. In summary, our contributions are three-fold:
  1. 1.
    General off-policy correction. We propose ALP, which unifies staleness of training policies and training–inference mismatch into a single ratio of the updated and inference policies. ALP is simple and efficient to implement and avoids the over-truncation and multi-threshold tuning required by prior two-ratio approaches.
  2. 2.
    Theory: smoother geometry and bounded discrepancy. We prove that with adaptive layerwise perturbation, KL divergence between the updated and inference distribution is bounded when $\sigma^{2}$ matches or exceeds the norm of the bias for the inference distribution from the training engine. This increases the probability that policy updates stay within a trust region and yields stable improvement. We further connect perturbation to loss smoothing, mitigating attraction to sharp, brittle optima.
  3. 3.
    Experiments: stability, performance, and ablations. Across single-turn and multi-turn agentic reasoning tasks, ALP improves stability and consistently outperforms MIS and Bypass. We further show that ALP reduces off-policy mismatch and exhibits more stable optimization dynamics, consistent with improved exploration and robustness. Ablations further reveal that perturbing all layers is most effective, and in partial settings, perturbations closer to lower layers tend to perform better than those restricted to upper layers.
2 Related Works
Off-policy
Off-policy optimization is a challenging setting in reinforcement learning. Trust Region Policy Optimization (TRPO) (Schulman et al., 2015) shows that when the KL-divergence is updated and the behavior policy is bounded, policy improvement is guaranteed. PPO (Schulman et al., 2017) further simplifies the optimization by proposing a surrogate clipping method. The off-policy issue becomes more severe in LLM reasoning tasks, especially in long CoT and multi-turn settings. To trade-off the bias and balance for importance ratios, sequence-level importance ratios (Zheng et al., 2025) and more stringent masking (Li et al., 2025) are developed to stabilize training. Moreover, the quantization and batching issues in the inference engine make the off-policy unavoidable. Recent works (Yao et al., 2025; Jiacai et al., 2025) correct the distribution by multiplying another importance ratio $\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}/\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}$ and clipping or masking the outliers.
Recent systems efforts aim to eliminate training-inference mismatch by enforcing bitwise-consistent inference, enabling closer-to on-policy RL in practice (He & Lab, 2025) or using higher precisions (Qi et al., 2025). These directions are promising, but maintaining strict bitwise consistency is computationally inefficient, and can be challenging under common production constraints such as changing execution engines, optional rollout quantization, fully asynchronous rollouts, and rapidly evolving kernels. Our work is complementary: rather than assuming mismatch can be removed everywhere, we propose an algorithmic correction that remains robust when such mismatch persists, providing a unified importance-ratio formulation that tolerates heterogeneous sources of off-policy deviation. Although those methods are effective, they only focus on a specific problem and current RL training needs to combine the techniques and tune the parameters separately. The algorithm becomes even more complicated in fully-asynchronous settings (Fu et al., 2025) since there are three importance ratios. Hence, our work develops a unified importance ratio instead of dividing the importance ratio into several parts to solve general off-policy problems.
Perturbation
Perturbation-based training is often motivated as a form of local averaging (randomized smoothing): optimizing the expected objective under small input/latent disturbances can reduce sensitivity to sharp, brittle regions of the loss landscape and improve robustness under distribution shift. This perspective has been used across a range of learning settings to stabilize optimization and enhance generalization.
A line of work has focused on enhancing model smoothness via perturbation-based sampling across diverse settings. For instance, Moreno-Barea et al. (2018); Yu et al. (2023) inject Gaussian noise to augment the training data and improve generalization. Certified robustness methods (Cohen et al., 2019; Salman et al., 2019; Lecuyer et al., 2019; Yang et al., 2020) pursue stability against adversarial attacks by introducing perturbations drawn from specific distributions during training. Beyond robustness, Li (2022); Pereira et al. (2021) demonstrate that perturbation injection can also improve performance across a range of tasks. Another related line of research focuses on diffusion models (Ho et al., 2020; Song et al., 2020; Dhariwal & Nichol, 2021; Saharia et al., 2022; Rombach et al., 2022), which model sample distributions as convolutions of Gaussian distributions and introduce Gaussian noise to the data during training. More recently, Shen & Meinshausen (2023); Hao et al. (2025) propose learnable perturbation mechanisms that adaptively enhance model performance. Related ideas also appear in diffusion models, where additional perturbations can mitigate train–test mismatch in iterative generation: Ning et al. (2023) reduce exposure bias by perturbing training inputs to better match the distribution encountered during sampling, while Li et al. (2023) alleviate exposure bias via sampling with shifted time steps.
3 Adaptive Layerwise Perturbation
3.1 Prior Approaches
For prompts $x\sim d_{0}$, the response $a$ is generated from the rollout policy $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(\cdot|x)$, where $\theta_{\mathrm{old}}$ is the parameter from the last step. We then update $\theta$ using training probability $\pi_{\theta}^{\mathrm{train}}$. Then, the standard GRPO
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=$ $\mathbb{E}_{x\sim d_{0},\{a_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(\cdot|x)}\bigg[\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{|a_{i}|}\min\Big\{\rho_{i,t}(\theta)A_{i},\text{clip}\left(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\right)A_{i}\Big\}\bigg],$ (1)
where $|a_{i}|$ denotes the length of $a_{i}$, $A_{i}$ is the group advantage, and $\rho_{t}(\theta)$ is token-level importance ratio at $t$-th token
$\rho_{i,t}(\theta)=\frac{\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
Sequence-level objective.
However, this token-level objective is biased when the difference between $\pi^{\mathrm{train}}_{\theta}$ and $\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}$ cannot be neglected. We can use the unbiased sequence level objective (Zheng et al., 2025),
$\rho_{i}(\theta)=\frac{\pi^{\mathrm{train}}_{\theta}(a_{i}|x)}{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i}|x)}=\prod_{t=1}^{|a_{i}|}\frac{\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
To balance the bias and variance, Group Sequence Policy Optimization (GSPO) (Zheng et al., 2025) proposes to use geometrical mean:
$\left(\frac{\pi^{\mathrm{train}}_{\theta}(a_{i}|x)}{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i}|x)}\right)^{\frac{1}{|a_{i}|}}.$
Masked importance sampling (MIS).
Since $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ and $\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}$ have a mismatch, to make algorithms robust to training-inference mismatch, a standard approach is to multiply by a correcting ratio (Yao et al., 2025; Liu et al., 2025):
$\mathcal{J}_{\mathrm{MIS}}(\theta)=$ $\mathbb{E}_{x\sim d_{0},\{a_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(\cdot|x)}\bigg[\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{|a_{i}|}\mathrm{Mask}\left\{\rho^{\prime}_{i,t},C\right\}\cdot\min\Big\{\rho_{i,t}(\theta)A_{i},\text{clip}\left(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\right)A_{i}\Big\}\bigg],$
where $C$ is a threshold parameter, the gradient is masked when $\rho^{\prime}_{i,t}>C$, and the additional importance ratio
$\rho^{\prime}_{i,t}=\frac{\pi^{\mathrm{train}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
This ratio can also be calculated in sequence level $\rho^{\prime}_{i}=\prod_{t}\rho^{\prime}_{i,t}$.
Bypass fail to stabilize training.
A direct idea to unify the importance ratio is directly replacing $\rho_{i}(\theta)$ with
$\frac{\pi^{\mathrm{train}}_{\theta}(a_{i,t}|x,a_{i,<t})}{\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})}.$
This approach is a baseline in training-inference mismatch works (Yao et al., 2025; Jiacai et al., 2025), but it does not work effectively. Specifically, since the family class of $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ has system bias from the training policy $\pi_{\theta_{\mathrm{old}}}$, the policy gradients have noises and biases, and thus making the policy more unstable and brittle. Empirically, Bypass has blow up ratio tails during full training steps in Figure 1 (Right). Even with the same checkpoint and the same rollouts, the (log-)ratio envelope can flare up in the low-probability tail without additional mechanisms to control mismatch (Fig. 2 middle).
3.2 Our Approach
For layer $h\in[H]$ in the model, suppose the dimension of its input embedding is $d_{h}$, we add a Gaussian  [^2] We focus on Gaussian perturbations in experiments and analysis; extending to other zero-mean perturbations is an interesting direction. perturbation variable $\delta^{h}\sim\mathcal{N}(0,\sigma_{h}I_{d_{h}})$ to the updated training policy. Denoting the perturbation variable $\delta=[\delta^{1},\ldots,\delta^{H}]^{T}\sim\mathcal{N}(0,\sigma I_{d})$, where $\sigma$ is the std vector and $d=\sum_{h=1}^{H}d_{h}$, we define the perturbed policy
$\pi_{\theta,\sigma}(a|x,\delta)=\Pi_{t=1}^{|a|}\pi_{\theta,\sigma}(a_{t}|x,a_{<t},\delta_{t}),$
where $\delta_{t}$ matches the hidden-state tensor shape and is sampled independently across token positions. Then, the loss function is
$\mathcal{J}_{\mathrm{ALP}}(\theta,\sigma)=$ $\mathbb{E}_{x\sim d_{0},\{a_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(\cdot|x),\delta_{i}\sim\mathcal{N}(0,\sigma I_{d})}\bigg[\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{|a_{i}|}\min\Big\{\rho^{\mathrm{ALP}}_{i,t}(\theta)A_{i},\text{clip}\left(\rho^{\mathrm{ALP}}_{i,t}(\theta),1-\epsilon_{l},1+\epsilon_{h}\right)A_{i}\Big\}\bigg],$ (2)
where the token-level importance ratio
$\rho^{\mathrm{ALP}}_{i,t}=\frac{\pi_{\theta,\sigma}(a_{i,t}|x,a_{i,<t},\delta_{i,t})}{\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i,t}|x,a_{i,<t})},$
and the sequence-level importance ratio
$\rho^{\mathrm{ALP}}_{i}=\prod_{t=1}^{|a_{i}|}\rho^{\mathrm{ALP}}_{i,t}=\frac{\pi_{\theta,\sigma}(a_{i}|x,\delta_{i})}{\pi^{\mathrm{infer}}_{\theta_{\mathrm{old}}}(a_{i}|x)}$ (3)
can also be applied to the loss correspondingly. Note that only the training policy is perturbed not the inference one.
3.3 Analysis: Robustness of ALP
Figure 2: Benefits of perturbation on smoothness: (Left) Toy simulation to show how perturbation smooths a sharp, spiky objective into a flatter surrogate, reducing sensitivity to local sharp maxima and promoting progress toward broader optima. (Middle, Right) Multi-turn, one-iteration controlled comparison. Starting from the same checkpoint and same batch of rollout samples, we perform $16$ off-policy updating steps without perturbation (middle) and with perturbation (right). Perturbation markedly shrinks the mismatch envelope, especially for low-probability tokens, thereby reducing extreme training–inference deviations.
The benefits of ALP can be stated in two aspects: both training-inference mismatch and smoothness. For simplicity, we conduct the theoretical analysis on one-layer perturbed policy. For each layer the perturbation $\delta$ is added to the embedding feature $x$, then ALP aims to learn a policy model with the expression as follows:
$\widetilde{\pi}_{\theta}(a|x)=\mathbb{E}_{\delta}\pi_{\theta}(a|x+\delta),$
For analytical convenience, we take the expectation over $\delta$. In practice, however, the algorithm samples $\delta$ only once per updating. This introduces a small discrepancy between the theoretical expectation and the single-sample approximation. We analyze the expectation form for clarity; in practice we use a single-sample approximation, which we find empirically sufficient in our setting.
Shrink the training-inference mismatch.
In the one-layer model, the training-inference mismatch can be formulated as
$\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\approx\pi_{\theta_{\mathrm{old}}}(a|x+\zeta),$
where the mismatch $\zeta$ is a zero-mean random variable.
For the original non-perturb method, taking the second-order Taylor expansion on KL distance, we have
$\quad\mathbb{E}_{x\sim d_{0}}\mathrm{KL}\left(\pi_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\approx\frac{1}{2}\zeta^{T}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}}\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)^{T}\zeta,$
which implies that
$\mathbb{E}_{\zeta}\mathbb{E}_{x\sim d_{0}}\mathrm{KL}\left(\pi_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\approx\frac{1}{2}\mathrm{Tr}\left(\mathbb{E}\zeta\zeta^{T}\cdot\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}}\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)^{T}\right).$ (4)
This approximation holds since the mismatch bias $\zeta$ is sufficiently small. Even with the small $\mathbb{E}\|\zeta\|_{2}^{2}$, as the eigenvalues of $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}}\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\pi_{\theta_{\mathrm{old}}}(a|x)^{T}$ can not be restricted during training process, a small training-inference mismatch is not guaranteed.
Now in ALP, with a proper perturbation variance $\sigma^{2}$, we could control the KL distance above effectively
Theorem 1 (Informal).
When perturbation is not too large to distort the original distribution, we have
$\mathbb{E}_{x\sim d_{0},\zeta}\mathrm{KL}\left(\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\leq\mathcal{O}\left(\frac{d\mathbb{E}\|\zeta\|_{2}^{2}}{2\sigma^{2}}\right),$ (5)
where $\mathcal{O}(\cdot)$ hides absolute constants and lower-order terms.
This theorem shows a trade-off on the scale of perturbation. Only when the perturbation std is small enough to stay close to the original distribution and large enough to cover the system noise norm $\|\zeta\|$, the KL-divergence between train and inference policy can be controlled effectively. The formal theorem is deferred to Theorem 3.
Moreover, TRPO shows the training is stabilized only when the distance between behavior and updated policy $\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\pi^{\mathrm{train}}_{\theta})$ is upper bounded by $\epsilon$ and the policy improvement is negatively affected by $\epsilon$ (Schulman et al., 2015). The distance is upper bounded by
$\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\widetilde{\pi}_{\theta})\leq\underbrace{\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\widetilde{\pi}_{\theta_{\mathrm{old}}})}_{\mathrm{Mismatch~Term}}+\underbrace{\mathrm{D}_{TV}(\widetilde{\pi}_{\theta_{\mathrm{old}}},\widetilde{\pi}_{\theta})}_{\mathrm{Controllable~Term}},$
where the first term is related to training-inference mismatch and the second term pertains to the training engine that is controllable and degrades to 0 when each batch has only one update. Hence, the instability mainly comes from Mismatch Term.
Without perturbation, as the mismatch between $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ can not be controlled from Equation (4), the condition $\mathrm{D}_{TV}(\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}},\pi_{\theta})\leq\epsilon^{\prime}$ in TRPO can not be easily satisfied. However, in ALP, Mismatch Term can be controlled by Theorem 3. Therefore, the updated policy $\widetilde{\pi}_{\theta}$ is more likely to fall into the trust region and guarantee training stability. In Figure 2 (Middle/Right), the quantile envelope shrinkage further validates this theory.
Improvement on Loss Smoothness.
On another side, adding perturbation can benefit the robustness of policy distribution, which implies that the smoothness can be smaller than the original algorithm. Without loss of generality, we neglect clipping and consider the standard reinforce loss and the perturbation loss:
$J(\theta)$ $=\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\left[\frac{\pi^{\mathrm{train}}_{\theta}(a|x)}{\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\cdot A(a)\right],$
$J_{\mathrm{ALP}}(\theta,\sigma)$ $=\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x),\delta\sim N(0,I)}\left[\frac{\pi^{\mathrm{train}}_{\theta}(a|x+\delta)}{\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\cdot A(a)\right].$
The main result is as follows (See the detailed proofs in Appendix C):
Theorem 2.
Define $\mathcal{I}(x,\theta)=\|\nabla^{2}_{\theta}J(\theta)\|_{2},\quad\widetilde{\mathcal{I}}(x,\theta)=\|\nabla^{2}_{\theta}J_{\mathrm{ALP}}(\theta,\sigma)\|_{2}$. If there exits a smallest constant $0<c<1$ such that
$\int\bm{1}\left(\mathcal{I}(x,\theta)\geq c/2\cdot\max_{x}\mathcal{I}(x,\theta)\right)dx<+\infty,$
with some proper distribution $\delta\sim\mathcal{N}(0,\sigma^{2}I)$, we have
$\sup_{x}\widetilde{\mathcal{I}}(x,\theta)\leq c\cdot\sup_{x}\mathcal{I}(x,\theta).$
This result demonstrate that when the loss function is unsmooth, i.e., the curvature measured by the Hessian spectral norm $\mathcal{I}(x,\theta)$ varies substantially across different input points $x$, ALP can achieve better smoothness than the original non-perturbed method.
Empirical analysis.
The theories are demonstrated by empirical diagnostics. First, as shown in Figure 2 (Left), by introducing Gaussian perturbations, these high-curvature regions are effectively smoothed out. Thus, perturbation transforms a sharp objective into a locally smoothed surrogate, preventing optimization from collapsing onto spiky maxima and favoring flatter regions that are less sensitive to small distribution shifts. Additionally, in a controlled one-step intervention from the same checkpoint (Figure 2, middle/right), adding perturbation substantially narrows the conditional quantile envelope of the log-ratio, especially for low-probability tokens that dominate tail risk. This indicates that ALP also suppresses extreme training-inference deviations in a single update step, increasing the likelihood that updates remain within a trust region around the rollout distribution (Schulman et al., 2015).
4 Experiments
We conduct experiments on single-turn reasoning and multi-turn agentic tasks respectively.
4.1 Single-turn Reasoning
4.1.1 Experimental settings
In the single-turn reasoning, we study mathematical reasoning tasks without tool-calling.
Dataset.
We construct the training dataset by merging the math subset from Guru RL-92k [^3] https://huggingface.co/datasets/LLM360/guru-RL-92k(Cheng et al., 2025) and $75k$ OpenR1444https://huggingface.co/datasets/weqweasdas/from_default_filtered_openr1_with_scores(Hugging Face, 2025) subset. Guru contains $91.9K$ high-quality samples spanning six diverse reasoning-intensive domains, processed through a comprehensive five-stage curation pipeline to ensure both domain diversity and reward verifiability. OpenR1 consists of 220k math problems with two to four reasoning traces generated by DeepSeek R1 for problems from NuminaMath 1.5. We employ the Math-Verify tool (Kydlíček, ) for automatic solution correctness verification. The details are deferred to Appendix D.
Training.
The implementations are based on the verl framework (Sheng et al., 2024), and we follow most of the parameter settings in verl. Detailedly, we apply the AdamW optimizer with learning rate $1\times 10^{-6}$ and adopt the clip higher trick (Yu et al., 2025) to clip the sampling ratio to an asymmetric range $(1-\epsilon_{l},1+\epsilon_{h})$. We conduct Seq-ALP (sequence-level ALP Equation (3)) and Token-ALP (token-level ALP Equation (2)) under the same training pipeline. The training is initialized from Qwen2.5-Math-1.5B-base (Yang et al., 2024). In each iteration, we sample $512$ prompts and rollout $n=8$ samples with temperature $1.0$ for each prompt, and the policy update number is $16$ in each iteration.
Benchmarks.
We evaluate the models’ performance on $5$ benchmarks: Math500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (He et al., 2024), AIME2024555https://huggingface.co/datasets/math-ai/aime24 and AIME2025666https://huggingface.co/datasets/math-ai/aime25. We use average@$32$ for evaluation, i.e., the accuracy is averaged over $32$ responses per prompt under temperature $1.0$. The models are allowed to generate $4096$ tokens. We compare with four baselines: Seq-Bypass, Seq-MIS, Token-MIS and vanilla GRPO that differ in how training-inference importance ratio $\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ and policy update importance ratio $\pi_{\theta}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}$ are formed and clipped. As provided in Table 1, in each row, token/seq means the aggregation way of the ratio is in token-level or sequence level. The definitions of those notations are provided in Sec. 3.1.
Baselines $\pi_{\theta}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}$ $\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ $\pi_{\theta}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$
GRPO token None None
Token-MIS token seq None
Seq-MIS seq seq None
Seq-Bypass None None seq
Table 1: Settings of parameters of importance ratios
4.1.2 Main Results
Method Math500 Minerva Math Olympiad Bench AIME24 AIME25 Average
GRPO 75.91 36.43 38.82 16.77 10.94 35.77
Seq-Bypass 76.21 35.23 38.52 16.35 7.81 34.82
Seq-MIS 77.08 36.75 39.06 15.21 9.58 35.54
Token-MIS 77.84 35.94 40.06 17.40 10.83 36.41
Token-ALP 78.10 37.27 40.77 21.46 11.77 37.87
Seq-ALP 77.84 37.06 40.28 16.98 11.98 36.83
Table 2: Single-turn evaluation (average@32, temperature $1.0$) across five math benchmarks. Token-ALP attains the highest average score.
Figure 3: Single-turn training dynamics under training-inference mismatch. We report step-wise (policy-update steps) optimization diagnostics (batch-averaged reward, gradient norm, policy entropy, and KL divergence between rollout and training policies) for GRPO, MIS, Bypass, and ALP. The reward mean is smoothed with a 10-step moving average.
ALP-type delivers the top performance.
Under single-turn evaluation in Table 2 across five math benchmarks, Token-ALP attains the highest overall average score ($37.87$), with Seq-ALP ranking second ($36.83$). Token-ALP leads on four out of five benchmarks (Math500, Minerva Math, Olympiad Bench, and AIME24), while Seq-ALP achieves the best result on AIME25 ($11.98$). In comparison, Token-MIS is the strongest baseline, while Seq-MIS and GRPO are slightly weaker overall ($35.54$ and $35.77$). Seq-Bypass performs the worst (Average $34.82$).
ALP consistently improves training stability.
The training dynamics in Figure 3 further explain these performance differences under training-inference mismatch. First, both token-level and seq-level ALP exhibit stable optimization throughout training: the gradient norm remains well-controlled (no sharp spikes), the policy entropy stays in a reasonable band (avoiding both collapse and runaway growth), and critically the training-inference KL remains near zero, indicating minimal drift between the rollout policy used for data collection and the policy effectively optimized during training. Second, methods without perturbation show clear instability signatures. In particular, GRPO and Seq-MIS undergo abrupt transitions late in training, characterized by large gradient-norm spikes, a pronounced training-inference KL surge, and a simultaneous entropy anomaly, which coincide with a sharp drop in reward mean – a hallmark of training collapse under mismatch. By contrast, Seq-Bypass and Token-MIS appear more stable than GRPO/Seq-MIS, but they still fall short of ALP in combining strong stability with the best overall accuracy: Token-MIS remains the strongest baseline, yet both Token-ALP and Seq-ALP provide consistently better aggregate performance while preserving stable training dynamics.
5 Multi-Turn Agentic Reasoning
5.1 Experimental settings
We choose multi-turn Tool-Integrated Reasoning (TIR) task with a Python interpreter to perform mathematical reasoning, where the agent can code to validate results, solve equations and so on. We adapt the codebase from SimpleTIR (Xue et al., 2025) and mask void turns (turns without tool calls and without final answers) in the loss. For the dataset, we merge the dataset Math3-5 from SimpleRL (Zeng et al., 2025), Deepscaler (Luo et al., 2025), and rStar2-Agent (Shang et al., 2025). ALP follows zero-RL setting and starts from Qwen2.5-7B-base. During training, the rollout batch size is $512$, and the mini update size is $32$. The maximum response length is set as 1616K, with a maximum of $5$ turns of code execution. The baselines and test benchmarks are the same with the single turn setting.
5.2 Main Results
Algorithm Math500 Minerva Math Olympiad Bench AIME24 AIME25 Average
GRPO 80.90 42.13 49.59 35.42 24.79 46.57
Seq-Bypass 81.74 39.21 51.08 37.19 24.06 46.66
Token-MIS 83.29 41.45 50.65 39.48 28.85 48.74
Seq-MIS 80.61 42.05 48.38 39.48 24.17 46.94
Token-ALP 83.48 43.18 51.55 38.65 31.25 49.62
Seq-ALP 84.29 43.10 52.75 43.85 28.65 50.53
Table 3: TIR evaluation: the test accuracy average@32 under temperature $1.0$ across the five benchmarks. ALP has the best performance.
Figure 4: Training dynamics: The leftmost plot illustrates that Seq-ALP maintains higher entropy than MIS baselines, ensuring sufficient exploration, while avoiding the unstable entropy growth ("blow-up") observed in Seq-Bypass. The center and right plots show that Seq-ALP maintains consistent convergence in both KL metrics. In contrast, Token-MIS exhibits significant instability in Train-Inference KL, and Seq-MIS suffers from sharp spikes in Policy Update KL, highlighting the robustness of the ALP method.
Figure 5: Pass@k performance analysis on AIME 2024 and AIME 2025 datasets for TIR tasks. The results compare ALP against baseline methods across varying rollout numbers ($k$). ALP consistently achieves the highest accuracy in the range of $k=16\sim 256$, indicating that adaptive latent perturbation significantly enhances the model’s exploration efficiency and solution diversity.
ALP has the best performance.
On multi-turn TIR benchmarks in Table 3, Seq-ALP achieves the highest overall average accuracy ($50.53$), improving over Token-MIS ($48.74$), Seq-MIS ($46.94$), and Seq-Bypass ($46.66$). The gains are broad-based: Seq-ALP is best on Math500 ($84.29$), Minerva Math ($43.10$), Olympiad Bench ($52.75$), and AIME24 ($43.85$), and is competitive on AIME25 ($28.65$) (slightly below Token-MIS at $28.85$), yielding the strongest aggregate performance across tasks. In contrast, although the accuracy of token and sequence level MIS does not collapse, their training-inference KL or policy-update KL blow up during the training.
ALP enhances training stability.
Figure 4 further show that Seq-ALP exhibits consistently well-behaved optimization dynamics. In particular, Seq-ALP maintains controlled train–inference KL without the pronounced excursions observed in Token-MIS, and avoids the sharp spikes in policy-update KL that characterize Seq-MIS. In addition, Seq-ALP maintains moderately higher entropy (supporting exploration) while avoiding the entropy blow-up in Seq-Bypass, indicating that ALP stabilizes both policy drift and update magnitude during training.
Remark 1 (Token level and sequence level).
An observation is that token-level ALP performs best in single-turn settings, while sequence-level ALP performs best in multi-turn settings. A plausible explanation is an efficiency–stability trade-off. Token-level ratios can be biased under policy drift, but in relatively stable single-turn training this bias is small, and clipping only affects the few tokens whose ratios exceed the threshold, thus preserving most learning signal. In contrast, sequence-level ratios are less biased in principle, but are more sensitive to outliers: extreme tokens can cause the entire trajectory to be clipped, discarding substantial amount of signal. As drift and mismatch become more severe in multi-turn training, the token-level bias can grow, making the lower-bias sequence-level variant more stable.
Perturbation promotes exploration efficiency.
Finally, the Pass@$k$ curves on AIME24 and 25 provide direct evidence that latent-space perturbations translate into more effective exploration (Figure 5). Across moderate-to-large rollout budgets $k$ (1625616–256), ALP consistently attains the highest Pass@$k$, with the advantage becoming more pronounced as $k$ increases, suggesting that perturbing hidden states helps generate a more diverse set of candidate solution trajectories and thus improves the probability of sampling a correct reasoning path under fixed decoding temperature.
6 Ablations
6.1 Perturbation targets
Figure 6: Perturbation-layer ablation in the single-turn setting. The results reveal a clear performance hierarchy: All layers >> Partial layers >> Logits only; Bottom layers >> Top layers.
Figure 7: Layer-wise perturbation diagnostics across target choices: (left) mean adaptive noise scale $\sigma$, (middle) distribution of perturbation-induced probability differences $|\Delta p|$, and (right) tail behavior measured by the 75th percentile of $|\Delta p|$ over training steps.
We study how the location of Gaussian perturbations affects both performance and stability. Starting from the same base setup, we inject perturbations into (i) all transformer layers (0–27), (ii) early layers (0–4), (iii) middle layers (12–16), (iv) late layers (19–23), or (v) the output logits only (no hidden-state perturbation).
Main finding: All layers >> partial layers >> logits-only.
Across single-turn benchmarks (Figure 6), perturbing all layers consistently achieves the strongest average accuracy, while restricting perturbations to a narrow band of layers is weaker. Logits-only perturbation is the least effective. This pattern is consistent with the view that ALP benefits from representation-level family enlargement rather than output noise alone, since logits-only perturbation is substantially weaker.
Multi-turn Ablation.
Algorithm Math500 Minerva Math Olympiad Bench AIME24 AIME25 Average
Layers 0-5 83.58 43.62 49.88 38.85 25.31 48.25
Layers 12-17 82.19 41.70 51.69 39.69 27.29 48.51
Layers 23-27 82.77 43.97 49.60 40.10 26.87 48.66
Layers 0-27 84.29 43.10 52.75 43.85 28.65 50.53
Table 4: Perturbation-layer ablation in the multi-turn setting. All-layer perturbation outperforms partial-layer perturbations.
The same conclusion holds in multi-turn agentic evaluation (Table 4). All-layer perturbation attains the best overall score across benchmarks, whereas partial-layer perturbations trail behind. Since multi-turn rollouts compound distributional drift over time, this result indicates that perturbing broadly across depth is particularly effective at maintaining robustness under sequential feedback and tool-induced distribution shift.
Why do partial or late-layer perturbations under-perform?
Layer ablations suggest that perturbing all layers is the most reliable choice across settings; in single-turn tasks, earlier-layer perturbations can outperform later-layer ones. We interpret this through a distribution-family lens: injecting perturbations into hidden states enlarges the effective policy family $\{\pi_{\theta,\sigma}\}$, providing additional degrees of freedom to absorb inference-time distortions and increase support overlap with the rollout distribution. In contrast, restricting perturbations to a narrow late-layer band expands the family in a more limited and output-local way, which is less effective at matching the rollout distribution and more likely to concentrate changes near the decision boundary. Consistent with this view, our diagnostics show that all-layer perturbation achieves the smallest tail deviations in probability ratios, whereas late-layer-only perturbation exhibits heavier tails.
7 Conclusions and Future Work
We introduced Adaptive Layerwise Perturbation (ALP), a simple and unified approach for stabilizing off-policy RL in large language models. ALP injects learnable perturbations into transformer hidden states when computing the updated policy, and optimizes a single importance ratio between the perturbed training distribution and the rollout distribution. This perspective addresses both policy staleness (multiple mini-batch updates per rollout) and training–inference mismatch within a single formulation, while avoiding the complexity of separately truncating multiple ratios. Empirically, ALP consistently improves stability and performance on both single-turn and multi-turn agentic reasoning tasks. Ablations confirm that perturbing hidden states is essential, with perturbing all layers performing best overall and lower-layer perturbations typically preferred when only a subset is perturbed.
Several directions remain promising. First, we will evaluate ALP in fully asynchronous RL systems where a single trajectory is generated under several sequential policies; this setting exacerbates off-policy effects and provides a stringent test of ALP’s ability to control heavy-tailed ratios. Second, we will extend ALP to Mixture-of-Experts (MoE) models, where routing decisions and expert imbalance can amplify distribution shift and instability; perturbations may need to be coupled with routing-aware constraints or expert-specific scales. More broadly, it would be valuable to develop principled strategies for allocating perturbation across layers and experts, and to understand how perturbation interacts with truncation-based methods, quantization, and other systems-level sources of mismatch.
Appendix A Technical Lemmas
Lemma 1 (Stam’s inequality).
For any two independent random variables $(X,Y)$, we have
$I(X+Y)\preceq\min\{I(X),I(Y)\},$
where $I(\cdot)$ is defined as the Fisher information matrix of the corresponding random variable.
Lemma 2.
If we assume $\delta\sim\mathcal{N}(0,\sigma^{2}I_{d})$, and define $\widetilde{\pi}_{\theta}(a|x):=\mathbb{E}_{\delta}\pi(a|x+\delta)$, there will be
$\nabla_{x}\ln\widetilde{\pi}_{\theta}(a|x)=\frac{1}{\sigma^{2}}\mathbb{E}_{\delta\sim q_{\theta}(\cdot|a,x)}[\delta],$
where
$q_{\theta}(\delta|a,x)=\frac{\pi_{\theta}(a|x+\delta)\phi(\delta)}{\widetilde{\pi}_{\theta}(a|x)},\quad\phi(\delta)\sim\mathcal{N}(0,\sigma^{2}I_{d}).$
Proof.
First, we could calculate the gradient of $\widetilde{\pi}_{\theta}(a|x)$ with respect to $x$ as:
$\nabla_{x}\widetilde{\pi}_{\theta}(a|x)$ $:=\int\nabla_{x}\pi_{\theta}(a|x+\delta)\phi(\delta)d\delta$
$=\int\nabla_{\delta}\pi_{\theta}(a|x+\delta)\phi(\delta)d\delta$
$=-\int\pi_{\theta}(a|x+\delta)\nabla_{\delta}\phi(\delta)d\delta$
$=\int\frac{\delta}{\sigma^{2}}\pi_{\theta}(a|x+\delta)\phi(\delta)d\delta$
$=\frac{\widetilde{\pi}_{\theta}(a|x)}{\sigma^{2}}\int\delta\frac{\pi_{\theta}(a|x+\delta)\phi(\delta)}{\widetilde{\pi}_{\theta}(a|x)}d\delta$
$=\frac{\widetilde{\pi}_{\theta}(a|x)}{\sigma^{2}}\mathbb{E}_{\delta\sim q_{\theta}(\cdot|a,x)}\delta,$
where the third and fourth equality use integration by parts (Stein’s identity for a Gaussian) to move the derivative from the complex term onto the Gaussian density $\phi(\delta)$, whose gradient has a closed form and whose boundary term vanishes as $\|\delta\|\to\infty$.
Then take it into the expression of $\nabla_{x}\ln\widetilde{\pi}_{\theta}(a|x)$, we have
$\nabla_{x}\ln\widetilde{\pi}_{\theta}(a|x)=\frac{1}{\sigma^{2}}\mathbb{E}_{\delta\sim q_{\theta}(\cdot|a,x)}[\delta].$
Appendix B Formal Statement and Proof for Theorem 1
Condition 1.
For any prompt $x$ and action $a$, we have
$\|\mathbb{E}_{\delta\sim q(\cdot|a,x)}[\delta]\|_{2}^{2}\leq Cd\sigma^{2},$
where $C>0$ is a constant and $q_{\theta}(\cdot|a,x)$ is the posterior distribution of $\delta$ conditioned on $(a,x)$.
Condition 2.
There exist a constant $0<\alpha\leq 1$ such that for any $a$ and $x$, we have
$\alpha\cdot\widetilde{\pi}_{\theta}(a|x)\leq\pi_{\theta}(a|x).$
Condition 1 requires that the squared $\ell_{2}$ norm of the posterior mean of the perturbation is bounded by $\mathcal{O}(d\sigma^{2})$, where $d$ is the input dimension and $\sigma^{2}$ is the variance of the Gaussian prior $\delta\sim\mathcal{N}(0,\sigma^{2}I_{d})$. This scaling is natural and aligns with standard concentration behavior in high-dimensional Gaussian settings. Condition 2 ensures that the perturbed policy $\widetilde{\pi}_{\theta}$ is uniformly dominated by the original policy $\pi\theta$ up to a constant factor $\alpha$. Intuitively, this condition guarantees that the injected perturbation is well controlled and does not overly distort the action distribution, which could otherwise lead to instability during training. In the limiting case $\delta\to 0$, the condition is trivially satisfied with $\alpha=1$.
To be specific, we can obtain:
Theorem 3 (Formal Version of Theorem 1).
With Condition 1 and Condition 2, we have
$\mathbb{E}_{x\sim d_{0},\zeta}\mathrm{KL}\left(\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\leq-\ln\alpha+\frac{Cd\mathbb{E}\|\zeta\|_{2}^{2}}{2\sigma^{2}}.$ (6)
The result in Theorem 3 shows that with properly learned value of $\sigma$, the training-inference mismatch can be controlled effectively during the training process.
Proof.
In ALP, we could obtain the KL distance as:
$\quad\mathbb{E}_{\zeta}\mathbb{E}_{x\sim d_{0}}\mathrm{KL}\left(\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)$
$=\mathbb{E}_{\zeta}\mathbb{E}_{x\sim d_{0}}\sum_{a}\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\ln\left(\frac{\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)}{\pi_{\theta_{\mathrm{old}}}(a|x+\zeta)}\right)$
$\leq\mathbb{E}_{\zeta}\mathbb{E}_{x\sim d_{0}}\sum_{a}\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\ln\left(\frac{\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)}{\alpha\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x+\zeta)}\right)$
$\approx-\ln\alpha+\frac{1}{2}\mathbb{E}_{\zeta}\zeta^{T}\mathbb{E}_{x\sim d_{0},a\sim\widetilde{\pi}_{\theta_{\mathrm{old}}}(\cdot|x)}\nabla_{x}\ln\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\nabla_{x}\ln\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)^{T}\zeta$
$\leq-\ln\alpha+\frac{1}{2}\mathbb{E}\|\zeta\|_{2}^{2}\mathbb{E}_{x\sim d_{0},a\sim\widetilde{\pi}_{\theta_{\mathrm{old}}}(\cdot|x)}\left\|\nabla_{x}\ln\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\right\|_{2}^{2}$
$=-\ln\alpha+\frac{1}{2\sigma^{4}}\mathbb{E}\|\zeta\|_{2}^{2}\mathbb{E}_{x\sim d_{0},a\sim\widetilde{\pi}_{\theta_{\mathrm{old}}}(\cdot|x)}\left\|\mathbb{E}_{\delta\sim q_{\theta}(\cdot|a,x)}\delta\right\|_{2}^{2}$
$\leq-\ln\alpha+\frac{Cd\mathbb{E}\|\zeta\|_{2}^{2}}{2\sigma^{2}}.$
where the first inequality is from Condition 2, the equality on the sixth line is from Lemma 2, and the last inequality is from Condition 1. So we could bound the KL distance as
$\mathbb{E}_{x\sim d_{0},\zeta}\mathrm{KL}\left(\widetilde{\pi}_{\theta_{\mathrm{old}}}(a|x)\parallel\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)\right)\leq-\ln\alpha+\frac{Cd\mathbb{E}\|\zeta\|_{2}^{2}}{2\sigma^{2}}.$ (7)
Appendix C Proof for Theorem 2
To start with, we denote $b:=max_{x}\mathcal{I}(x,\theta)$, and there exists $c\in(0,2)$ such that
$\int\bm{1}(\mathcal{I}(x,\theta)\geq c/2\cdot b)\cdot dx=C<+\infty.$
Then for any $\epsilon>0$, we can choose proper std value $\sigma$, such that for any r satisfying that the volume of $r$-radius ball is no larger than $C$, i.e, $\mu\{\|\delta\|_{2}\leq r\}\leq C$, we have
$\mathcal{P}\left(\|\delta\|_{2}\leq r\right)\leq\epsilon.$ (8)
Then for any $x,a$, we have
$\quad\left\|\mathbb{E}_{\delta}\nabla^{2}_{\theta}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\left[\frac{\pi_{\theta}(a|x+\delta)}{\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\cdot\hat{A}(a)\right]\right\|_{2}\leq\mathbb{E}_{\delta}\left\|\nabla^{2}_{\theta}\mathbb{E}_{a\sim\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\left[\frac{\pi_{\theta}(a|x+\delta)}{\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}(a|x)}\cdot\hat{A}(a)\right]\right\|_{2}$
$=\int_{\mathcal{I}(x+\delta,\theta)\geq bc/2}\mathcal{I}(x+\delta,\theta)\phi(\delta)d\delta+\int_{\mathcal{I}(x+\delta,\theta)<bc/2}\mathcal{I}(x+\delta,\theta)\phi(\delta)d\delta$
$\leq b\cdot\int_{\mathcal{I}(x+\delta,\theta)\geq bc/2}\phi(\delta)d\delta+\frac{bc}{2}$
$\leq(\frac{c}{2}+\epsilon)\cdot b,$
where the last inequality is from the fact that the measure of set satisfying $\mathcal{I}(x,\theta)\geq c/2\cdot b$ equals to $C$, and for a Gaussian distribution, the probability on such set is smaller than the probability on a zero-center ball with volume $C$. Combing this fact with the assumption in Equation (8), we could derive the last inequality.
While $\epsilon\leq\frac{c}{2}$, we can finish the proof.
Appendix D Additional Expeimental Details
[…]
For the clipping or masking threshold parameters, they are chosen to achieve stable performance for different settings respectively. Specifically, for token-level importance ratio, the lower clipping ratio is $0.2$, and the higher clipping ratio is $0.2$ or $0.28$; for sequence-level importance ratio, the lower clipping ratio is $0.5$, and the higher one is $3.0$. Besides, for MIS, the masking threshold is $2.0$.
Figure 8: Single-turn evaluation on the test dataset. We compare final-task performance of GRPO, MIS, Bypass, and ALP, highlighting that stable optimization dynamics translate into improved reward/generalization.
D.1 Hyperparameter sensitivity
We study the sensitivity of ALP to two key hyperparameters: the initial Gaussian noise standard deviation (std) and the learning rate (lr) for the perturbation parameters. We evaluate the resulting policies on Math500, Minerva Math, OlympiadBench, AIME24, and AIME25, and report the average@32 score.
Figure 9: Hyperparameter sensitivity in the single-turn setting. We vary the initial Gaussian noise standard deviation (std) and its learning rate (lr), and report average@32 across Math500, Minerva Math, OlympiadBench, AIME24, and AIME25.
Figure 9 shows that performance curves are tightly clustered across a broad range of (std, lr) values and improve monotonically over training, suggesting that ALP is not overly sensitive to these hyperparameters. Based on the final-stage performance, we use std $1\times 10^{-4}$ and lr $5\times 10^{-4}$ as the default configuration in subsequent experiments.
Figure 10: Token-level training-inference mismatch at step 140 in the multi-turn setting. Each point is one response token, with x-axis denoting the training probability $\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}$ and y-axis denoting the rollout-time token probability $\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$. The red dashed diagonal is $y=x$; tighter concentration around this line indicates smaller training-inference mismatch. The Pearson correlation coefficient $r$ is reported in each panel title.